# Don’t Pay Attention

### v1

@@ 3 Experiments
3.1 Experimental Setup
We compare Avey against three leading open-source models, namely, Mamba [42], RWKV-7 [43], and Transformer++, extended to the strongest known architectural recipe of the standard Transformer [46] (see Appendix A for details). All models were trained using their best-known hyperparameters under a fixed budget of 100 billion tokens drawn from the FineWeb dataset [41]. Complete training and model hyperparameters for all the baselines are provided in Appendix A.
To assess each model’s accuracy, we employed a suite of widely used NLP benchmarks, including ARC-E and ARC-C [11], HellaSwag [105], PIQA [5], OBQA [61], SIQA [82], and Winogrande [80]. Additionally, we evaluated the long-context retrieval capabilities of all the models using the standard Single Needle-In-A-Haystack (S-NIAH) benchmark suite from RULER [38]. Full details of all the benchmarks and additional experimental setups are included in Appendix A.
3.2 Design Choices and Ablations
@@
RMSNorm or LayerNorm? RMSNorm Appendix H
LR: to decay or not to decay? Yes, cosine decay with peak LR of $1\text{e}{-3}$ Appendix I
Best values for sequence length $N$, split size $S$, and top-$k$ splits? $N=512$, $S=64$, $k=7$ Appendix G
Activation in the enricher? Yes, ReLU2 Appendix B
Activation in the contextualizer? No Appendix C
Deeper model and narrower embeddings, or shallower model and wider embeddings? Deeper model and narrower embeddings Appendix F
Weight ranked splits? Yes, using normalized scores Appendix K
Enrich embeddings before contextualization? Yes, by 4x Appendix K, D
@@
Bypass uncontextualized features to the fuser? Yes, 50% of each enriched embedding Appendix K, E
Static or dynamic parametrization for the contextualizer? Dynamic parametrization Appendix K
Replace the contextualizer with self-attention? No Appendix K
Besides enabling extrapolation past the context window, does the ranker improve performance? Yes Appendix K
We conducted over 200 experiments to investigate and ablate several key design choices in our architecture. These included, whether to: (1) utilize activation functions in the enricher and contextualizer; (2) trade off additional layers for reduced embedding dimensionality (or vice versa) under a fixed parameter budget; (3) apply weights to ranked splits; (4) enrich embeddings; (5) bypass uncontextualized features to the fuser; (6) use dynamic parameterization in the neural processor; and (7) replace the neural contextualizer with self-attention. Furthermore, we evaluated Avey without the ranker to assess its contribution to downstream task performance, beyond its primary role of enabling effective extrapolation past the trained context window. Table 1 summarizes our findings and provides references to the corresponding experiments that support each conclusion.
3.3 Short-Range Benchmark Results
Table 2: Zero-shot performance across multiple NLP tasks.
Model ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
Avey-153M 24.37 42.33 39.36 31.40 68.37 39.13 51.28 42.32
Transformer++-152M 23.63 43.17 39.32 29.80 67.01 38.89 50.22 41.72
Mamba-144M 24.17 43.53 40.55 30.40 68.32 39.41 52.72 42.73
RWKV-7-168M 24.17 43.01 41.55 29.67 68.72 39.17 51.09 42.48
Avey-496M 27.50 48.95 51.82 32.47 72.49 40.15 54.38 46.82
Transformer++-488M 26.73 48.09 52.66 31.73 72.13 39.93 55.25 46.65
Mamba-500M 28.64 51.02 54.15 34.47 73.03 40.84 55.49 48.23
RWKV-7-501M 27.13 49.37 54.54 36.27 73.58 39.40 55.72 48.00
Avey-1.52B 31.26 56.55 61.42 36.80 75.61 42.00 57.06 51.53
Transformer++-1.5B 30.00 56.29 63.87 38.00 76.01 42.24 61.38 52.54
Mamba-1.4B 32.13 57.74 63.74 36.85 76.19 42.00 60.40 52.72
RWKV-7-1.5B 32.94 59.05 64.43 37.13 76.84 41.71 60.06 53.17
We now evaluate Avey on standard autoregressive language modeling benchmarks, comparing it against Transformer++, Mamba, and RWKV-7 across three model sizes, small, medium, and large, as defined in Section 3.1. We utilize a range of widely used zero-shot downstream evaluation tasks, all detailed in Section 3.1. Table 2 summarizes the results. With small models, Avey, Mamba, and RWKV-7 outperformed Transformer++ by average margins of 1.43%, 2.41%, and 1.82%, respectively. Mamba and RWKV-7 slightly exceeded Avey’s performance, with average margins of 0.9% and 0.3%, respectively. With medium models, Avey, Mamba, and RWKV-7 again outperformed Transformer++ by averages of 0.3%, 3.4%, and 2.9%, respectively. Lastly, with large models, Avey underperformed Transformer++ by an average of 1.9%, while Mamba and RWKV-7 marginally outpaced it by 0.71% and 1.19%, respectively.
The results above assume a fixed training budget of 100B tokens. To better understand how Avey scales with increasing model size, we conducted additional experiments following the Chinchilla scaling laws [35], which recommend increasing the number of training tokens proportionally with model size. Consequently, we adjusted the number of training steps and tokens to align with these laws. Appendix J outlines the configurations of the trained models, including the numbers of layers, embedding dimensions, training steps, learning rates, and total training tokens. The methodology of these experiments closely follows that of [25], with slight modifications (e.g., to accommodate parameter budget constraints). As demonstrated in Appendix J, Avey exhibits more favorable scaling behavior than Transformer++, Mamba, and RWKV-7, particularly when both model size and token count are increased proportionally in accordance with the Chinchilla scaling laws.
Finally, to provide deeper insights into the behaviors of the compared models, we performed a detailed statistical analysis, presented in Appendix L. The analysis reports the standard deviation, standard error, and 95% confidence interval for each model, computed over its final three checkpoints—captured at 5-billion-token intervals (i.e., at 90B, 95B, and 100B tokens). The results revealed meaningful variance among models. For example, in the small model regime, although Avey does not surpass Mamba in mean performance, their confidence intervals overlap substantially within the range (42.06, 43.24), indicating that the two models are statistically comparable and that Avey may outperform Mamba in some runs. A similar, though narrower, overlap is observed between Avey and RWKV-7. In the large model regime, while Avey does not exceed Transformer++ in average performance, their confidence intervals also exhibit significant overlap, suggesting that Avey could potentially outperform Transformer++ in certain cases. A more limited overlap is seen between Avey and Mamba, implying that although Mamba generally performs better, Avey may outperform it in some runs.
3.4 Long-Range Benchmark Results
Figure 5: Performance comparison between Transformer++, Mamba, RWKV-7, and Avey on S-NIAH-1 and S-NIAH-2. The x-axis denotes the lengths of haystacks (i.e., documents with distractor texts, varying from 2k to 64k tokens). All models use 0.5B parameters. Similar results are shown in Appendix M for other model sizes.
In this section, we evaluate Avey, Transformer++, Mamba, and RWKV-7 on benchmarks designed to assess performance on tasks with long-range dependencies. Specifically, we use the standard Single Needle-In-A-Haystack (S-NIAH) benchmark suite from RULER[38], as described in Section3.1. The S-NIAH suite includes multiple variants, notably S-NIAH-1 (pass-key retrieval) and S-NIAH-2 (number in haystack). S-NIAH-1 involves retrieving the specific value associated with a given key (the pass-key) from a lengthy distractor text (the haystack) containing many key-value pairs. The pass-key, serving as the needle, appears only once, and the model must accurately recall its corresponding value regardless of its position in the haystack. S-NIAH-2 is similar to S-NIAH-1 but poses a greater challenge, whereby the value to be retrieved is numerical (e.g., a random 9-digit number). This task requires exact recall, where even a single-digit error is considered incorrect, thereby testing the model’s precision in extracting structured numerical information from long haystacks.
Fig. 5 demonstrates the results of Avey, Transformer++, Mamba, and RWKV-7 on both S-NIAH-1 and S-NIAH-2 benchmarks. As described in Section 3.1, Transformer++, Mamba, and RWKV-7 were all trained with a context window of 2,048 tokens. As shown, Transformer++ performs strongly on both benchmarks as long as the haystack length remains within its trained context window. Once the haystack’s length exceeds its context window width, Transformer++ fails to recall the correct values associated with the keys. In contrast, Mamba and RWKV-7 demonstrate some ability to generalize beyond their training windows, but their performance also declines significantly as the haystack length increases far beyond those limits. On the flip side, Avey achieves good performance across both benchmarks, despite being trained with a context window of only 512 tokens. For instance, on S-NIAH-2 with a 64k-token haystack, Avey outperforms Mamba and RWKV-7 by averages of 85.25% and 23.6%, respectively. In addition, on S-NIAH-1 under the same 64k-token setting, Avey achieves an accuracy of 97.8%, while Mamba and RWKV-7 drop to 0% and 0.8%, respectively.
Interestingly, Avey’s performance tends to improve as the haystack length increases, highlighting its strong extrapolative capability. This behavior can be attributed to the fact that as the haystack length (i.e., sequence length $N$) grows, the candidate pool from which the ranker selects the top-$k$ splits for contextualization also expands. As discussed in Appendix G, a larger $N$ enables the ranker to identify and retrieve more relevant splits while discarding less informative ones, thereby improving the overall quality of contextualization and potentially enhancing performance. This effect is further supported by the results in Appendix K, where the inclusion of the ranker led to measurable performance gains. Notably, embeddings containing a needle—whether in S-NIAH-1 or S-NIAH-2—are not processed in isolation but rather contextualized alongside other embeddings. As such, an improved quality of contextualization driven by the ranker may contribute to more accurate value recall. However, whether this mechanism fully explains Avey’s increasing performance with longer haystacks remains uncertain, and further interpretability studies are needed to better understand the underlying drivers of this behavior.
4 Related Work
Extensive research has focused on addressing the limitations of the Transformer and exploring alternative RNN-like architectures. Appendix O provides a comprehensive survey of these efforts and offers a comparison with Avey.
5 Conclusion
In this paper, we introduced Avey, a new foundational architecture for autoregressive language modeling. Unlike traditional models, Avey relies neither on recurrence nor attention. Instead, it employs a neural approach to enrich and contextualize embeddings. Additionally, it leverages a ranker that enables the model to flexibly and effectively handle sequences of arbitrary lengths, despite being trained with only a small context window. We hope this work lays the groundwork for future research and inspires further advances in scalable and effective language modeling.
Limitations
We define the scope of our work and discuss limitations in Appendix P.
Appendix A Experimental Methodology
In this section, we describe the experimental methodology employed throughout the paper. To begin with, we adopted a cascaded search process to identify the best configuration of Avey. Specifically, we began with a baseline version of the neural processor—excluding the ranker—and sequentially explored several architectural design choices. After each empirical conclusion regarding a specific architectural element, we integrated that element into the processor (one at a time) and resumed the search process from the updated configuration. This cascaded process is captured chronologically in Appendices B, C, D, E, and F.
To elaborate, we started with an expansion factor of 4×4$\times$4 × in the enricher, a tail size of 50% (i.e., half of each expanded embedding is forwarded to the contextualizer), RMSNorm [108] as a normalization technique, no activation functions in the enricher and contextualizer, global batch size of 0.5M, context width of 1024, and a constant learning rate of $1e{-3}$. As we empirically verified and decided upon each architectural element, we updated the processor accordingly. For example, after determining that ReLU2 was the most effective activation function for the enricher, we integrated it into the model and proceeded with the remaining search.
After finalizing the above exploratory set of experiments, we incorporated the ranker into the neural processor and conducted an extensive study—comprising over 138 training and inference runs—to identify the optimal sequence length (i.e., $N$), split size (i.e., $S$), and number of top-$k$ splits (i.e., $k$). The results of this sensitivity study are summarized in Appendix G. Following this, we evaluated the best normalization technique (Appendix H) as well as the optimal peak learning rate and learning rate schedule (Appendix I) for the full architecture.
All the experiments described above were conducted using a 145-million-parameter model trained on 10 billion tokens from the FineWeb dataset [^6] This dataset is released under the Open Data Commons Attribution License (ODC-By) v1.0. [41] (specifically, the sample-100BT subset of FineWeb). The results of these experiments informed the following final selection of training and model hyperparameters for Avey across three parameter scales, 153M (small), 496M (medium), and 1.52B (large).
Training and Model Hyperparameters of Avey:
  • Training hyperparameters:
    • Optimizer: AdamW
    • Betas: (0.9, 0.95)
    • Epsilon: $1\text{e}{-12}$
    • Peak learning rate: $1\text{e}{-3}$
    • Schedule: Cosine decay to 10% of the peak learning rate, with no warmup
    • Batch size: 0.5M for the small and medium models, and 1M for the large model
    • Gradient norm clip: 1.0
    • Weight decay: 0.1 (applied only to matrices)
  • Model hyperparameters:
    • All models:
      • Context width: 512
      • Split size ($S$): 64
      • Number of top-$k$ splits: 7
      • Vocabulary size: 50,304
      • Expansion factor: 4
      • Tail size: 0.5
    • Small model (153M parameters):
      • Embedding dimension: 768
      • Number of layers: 26
    • Medium model (496M parameters):
      • Embedding dimension: 768
      • Number of layers: 104
    • Large model (1.52B parameters):
      • Embedding dimension: 2048
      • Number of layers: 48
For baselines, we compared Avey against three leading open-source models, namely, Mamba [42], RWKV-7 [43], and Transformer++ [46]. For Transformer++, we implemented the strongest architectural recipe known to us, incorporating rotary positional encodings [89], SwiGLU MLPs [84], and RMSNorm in place of LayerNorm [108]. All models were trained using their best-known hyperparameters (see details below, assuming three model sizes, small, medium, and large) under a fixed budget of 100 billion tokens drawn from the aforementioned FineWeb dataset  [^7] More precisely, all models, including Avey, Transformer++, Mamba, and RWKV-7 were trained for 1 epoch over the sample-100BT subset of FineWeb.. For consistency and comparability, we used the p50k_base tokenizer [64] across all the models, as it aligns with the GPT-2-derived token counts reported for this dataset.
Training and Model Hyperparameters of Transformer++:
  • Training hyperparameters:
    • Optimizer: AdamW
    • Betas: (0.9, 0.95)
    • Epsilon: $1\text{e}{-12}$
    • Peak learning rates:
      • Small model: $3\text{e}{-3}$
      • Medium model: $1.5\text{e}{-3}$
      • Large model: $1.25\text{e}{-3}$
    • Schedule: A linear warmup for 10% of steps, followed by cosine decay to 10% of the peak learning rate
    • Batch size: 0.5M for the small and medium models, and 1M for the large model
    • Gradient norm clip: 1.0
    • Weight decay: 0.1 (only applied to matrices)
  • Model hyperparameters:
    • All models:
      • Context width: 2048
      • Vocabulary size: 50,304
      • Intermediate size in FFN: 4$\times$ the embedding dimension
    • Small model (152M parameters):
      • Embedding dimension: 768
      • Number of layers: 12
      • Number of heads: 12
    • Medium model (488M parameters):
      • Embedding dimension: 1024
      • Number of layers: 26
      • Number of heads: 16
    • Large model (1.5B parameters):
      • Embedding dimension: 1664
      • Number of layers: 32
      • Number of heads: 16
Training and Model Hyperparameters of Mamba:
  • Training hyperparameters:
    • Optimizer: AdamW
    • Betas: (0.9, 0.95)
    • Epsilon: $1\text{e}{-12}$
    • Peak learning rates:
      • Small model: $3\text{e}{-3}$
      • Medium model: $1.5\text{e}{-3}$
      • Large model: $1.0\text{e}{-3}$
    • Schedule: A linear warmup for 10% of steps, followed by cosine decay to 10% of the peak learning rate
    • Batch size: 0.5M for the small and medium models, and 1M for the large model
    • Gradient norm clip: 1.0
    • Weight decay: 0.1 (applied only to matrices)
  • Model hyperparameters:
    • All models:
      • Context width: 2048
      • Vocabulary size: 50,304
      • All hyperparameters other than the ones specified are left at their default values according to [42]
    • Small model (144M parameters):
      • Embedding dimension: 768
      • Number of layers: 28
    • Medium model (500M parameters):
      • Embedding dimension: 1280
      • Number of layers: 42
    • Large model (1.4B parameters):
      • Embedding dimension: 2048
      • Number of layers: 52
Training and Model Hyperparameters of RWKV-7:
  • Training hyperparameters:
    • Optimizer: AdamW
    • Betas: (0.9, 0.95)
    • Epsilon: $1\text{e}{-12}$
    • Peak learning rates:
      • Small model: $6\text{e}{-4}$
      • Medium model: $4\text{e}{-4}$
      • Large model: $4\text{e}{-4}$
    • Schedule: Cosine decay to 10% of the peak learning rate, with no warmup
    • Batch size: 1M for the small and medium models, and 2M for the large model
    • Gradient norm clip: 1.0
    • Weight decay: 0.1 (applied only to matrices)
  • Model hyperparameters:
    • All models:
      • Context width: 2048
      • Vocabulary size: 50,304
      • All hyperparameters other than the ones specified are left at their default values according to [43]
    • Small model (168M parameters):
      • Embedding dimension: 768
      • Number of layers: 12
    • Medium model (501M parameters):
      • Embedding dimension: 1024
      • Number of layers: 30
    • Large model (1.5B parameters):
      • Embedding dimension: 2048
      • Number of layers: 24
To compare all models, we employed a suite of widely used NLP benchmarks, including ARC-E and ARC-C (for scientific reasoning and reading comprehension) [11], HellaSwag (for commonsense inference) [105], PIQA (for physical reasoning) [5], OBQA (for open-book science reasoning) [61], SIQA (for social interaction understanding) [82], and Winogrande (for coreference and commonsense reasoning) [80]. In addition, we evaluated long-context retrieval capabilities using the standard Single Needle-In-A-Haystack (S-NIAH) benchmark suite from RULER [38], which measures a model’s ability to extract pass-keys from large distractor corpora, with sequence lengths ranging from 2k to 64k tokens. All evaluations were conducted using the widely adopted LM Evaluation Harness from EleutherAI [23], consistent with the prior work in the field.
For all models, we reported performance in terms of benchmark accuracy [^8] Additionally, throughout the paper, all reported perplexity values specifically refer to training perplexity.. Specifically, we used normalized accuracy (acc-norm) from the LM Evaluation Harness whenever available. For each model, the reported score on each NLP benchmark is the average accuracy across its final three checkpoints (taken at 90B, 95B, and 100B tokens) to account for variability due to training randomness. Complete benchmark results across these checkpoints, along with key summary statistics and discussions, are provided in Appendix L.
Finally, all training and evaluation runs were executed on 208 NVIDIA H200 GPUs, with mixed-precision (bfloat16) enabled for training. The total training time for all models—Avey, Transformer++, Mamba, and RWKV-7—across the three presented model sizes and 100B training tokens is estimated to be approximately 80–90 hours, assuming optimal parallelization across the 208 GPUs and using the implementations referenced above. To avoid potential sources of randomness and ensure consistency across results, we disabled Torch Compile during all design choice and sensitivity experiments. For the ablation studies and final training runs, however, Torch Compile was enabled whenever possible to accelerate training. Additionally, we fixed the random seed to 11 (arbitrarily chosen) for all training runs to further reduce variability due to stochastic effects.
Appendix B Activation or No Activation in the Enricher
Table 3: Avey’s performance without and with an activation function in the enricher. The study involves only the neural processor within Avey and trains it on 10B tokens.

### v2

@@ 3 Experiments
3.1 Methodology
We now describe our experimental methodology in detail. To begin with, we adopt a cascaded search to identify the best configuration of Avey. Starting from a baseline neural processor (excluding the ranker), we sequentially evaluate individual architectural choices. After each empirical finding, we integrate the best-performing architectural element into the processor (one at a time) and resume the search process from the updated configuration. This cascaded procedure unfolds chronologically in Sections 3.2.1, 3.2.2, 3.2.3, 3.2.4, and 3.2.5.
More precisely, we initiate the cascaded search with an expansion factor of $4\times$ in the enricher, a tail size of $50\%$ (i.e., half of each expanded embedding is forwarded to the contextualizer), RMSNorm (zhang2019rmsnorm) for normalization, no activations in either the enricher or contextualizer, a global batch size of 0.5M, a context width of 1024, and a constant learning rate of 1e31e{-3}. As each architectural element is explored and empirically selected, we incorporate it into the processor and resume the search from the updated configuration. For instance, after identifying $\mathrm{ReLU}^{2}$ as the most effective activation for the enricher, we included it in the architecture and proceeded with the remaining explorations.
After completing the exploratory experiments described above, we integrated the ranker into the neural processor and conducted an extensive study—spanning over 138 training and inference runs—to identify the optimal sequence length ($N$), split size ($S$), and number of top-$k$ splits ($k$). The results of this sensitivity analysis are summarized in Section 3.2.6. Subsequently, we evaluated normalization strategies in Section 3.2.7 and tuned the peak learning rate and learning-rate schedule for the full architecture in Section 3.2.8.
Table 1: Training and model hyperparameters for Avey, Transformer++, Mamba, and RWKV-7 at three scales.
Avey Transformer++ Mamba RWKV-7 Scale Small Medium Large Small Medium Large Small Medium Large Small Medium Large # of Parameters 153M 496M 1.52B 152M 488M 1.5B 144M 500M 1.4B 168M 501M 1.5B Optimizer AdamW AdamW AdamW AdamW Betas (0.9, 0.95) (0.9, 0.95) (0.9, 0.95) (0.9, 0.95) Epsilon $1\mathrm{e}{-12}$ $1\mathrm{e}{-12}$ $1\mathrm{e}{-12}$ $1\mathrm{e}{-12}$ Peak learning rate $1\mathrm{e}{-3}$ $3\mathrm{e}{-3}$ $1.5\mathrm{e}{-3}$ $1.25\mathrm{e}{-3}$ $3\mathrm{e}{-3}$ $1.5\mathrm{e}{-3}$ $1.0\mathrm{e}{-3}$ $6\mathrm{e}{-4}$ $4\mathrm{e}{-4}$ $6\text{e}{-4}$ $4\text{e}{-4}$ $4\text{e}{-4}$ Schedule Cosine decay to 10% of peak; no warmup Linear warmup 10% of steps; cosine decay to 10% Linear warmup 10% of steps; cosine decay to 10% Cosine decay to 10% of peak; no warmup Batch size 0.5M 0.5M 1M 0.5M 0.5M 1M 0.5M 0.5M 1M 1M 1M 2M Gradient norm clip 1.0 1.0 1.0 1.0 Weight decay 0.1 (matrices only) 0.1 (matrices only) 0.1 (matrices only) 0.1 (matrices only) Context width 512 2048 2048 2048 Split size ($S$) 64 N/A N/A N/A Top-$k$ splits 7 N/A N/A N/A Vocabulary size 50,304 50,304 50,304 50,304 Expansion factor 4 N/A N/A N/A Tail size 0.5 N/A N/A N/A Embedding dimension 768 768 2048 768 1024 1664 768 1280 2048 768 1024 2048 Number of layers 26 104 48 12 26 32 28 42 52 12 30 24 Number of heads N/A 12 N/A 16
All the aforementioned experiments were conducted with a 145M-parameter model trained on 10B tokens from the FineWeb dataset [^8] This dataset is released under the Open Data Commons Attribution License (ODC-By) v1.0. (fineweb2023) (specifically, the sample-100BT subset). These results informed the final choices of training and model hyperparameters for Avey across three parameter scales, 153M (small), 496M (medium), and 1.52B (large), as reported in Table 1.
For baselines, we compare Avey against three leading open-source models, namely, Transformer++ (nanogpt2023), Mamba (mambai2023), and RWKV-7 (rwkvi2023). For Transformer++, we implement the strongest recipe known to us, incorporating rotary positional encodings (RoPE) (su2024roformer), SwiGLU MLPs (shazeer2020glu), and RMSNorm in place of LayerNorm (zhang2019rmsnorm). All models are trained with their best-known hyperparameters (see Table 1) at three parameter scales (i.e., small, medium, and large) under a fixed budget of 100B tokens drawn from the FineWeb dataset [^9] More precisely, all models—Avey, Transformer++, Mamba, and RWKV-7—are trained for one epoch over the sample-100BT subset of FineWeb.. For consistency and comparability, we use the p50k_base tokenizer (tiktoken2022) across all models, matching the GPT-2–derived token counts reported for this dataset.
To compare all models, we employ a suite of widely used NLP benchmarks, including ARC-E and ARC-C (for scientific reasoning and reading comprehension) (clark2018think), HellaSwag (for commonsense inference) (zellers2019hellaswag), PIQA (for physical reasoning) (bisk2020piqa), OBQA (for open-book science reasoning) (mihaylov2018can), SIQA (for social interaction understanding) (sap2019socialiqa), and Winogrande (for coreference and commonsense reasoning) (sakaguchi2021winogrande). In addition, we evaluate long-context retrieval capabilities using the standard Single Needle-In-A-Haystack (S-NIAH) benchmark suite from RULER (hsieh2024ruler), which measures a model’s ability to recover passkeys from large distractor contexts at sequence lengths ranging from 2k to 64k tokens. All evaluations are conducted with the widely adopted LM Evaluation Harness from EleutherAI (gao2021framework), consistent with prior work (gu2023mamba).
For all models, we report performance in terms of benchmark accuracy [^10] Additionally, all reported perplexity values in this paper refer to training perplexity.. When available, we use the normalized-accuracy metric (acc_norm) from the LM Evaluation Harness. To mitigate variance due to training stochasticity, each model’s score on a given NLP benchmark is computed as the mean accuracy over its final three checkpoints (at 90B, 95B, and 100B tokens). Complete results across these checkpoints, along with summary statistics and discussion, are provided in Appendix B.
@@ 3.1 Methodology
Activation in the enricher? Yes, ReLU2 Section 3.2.1
Activation in the contextualizer? No Section 3.2.2
Enrich embeddings before contextualization? Yes, by 4x Sections 3.3, 3.2.3
@@
Bypass uncontextualized features to the fuser? Yes, 50% of each enriched embedding Sections 3.3, 3.2.4
Deeper model and narrower embeddings, or shallower model and wider embeddings? Deeper model and narrower embeddings Section 3.2.5
Best values for sequence length $N$, split size $S$, and top-$k$ splits? $N=512$, $S=64$, $k=7$ Section 3.2.6
RMSNorm or LayerNorm? RMSNorm Section 3.2.7
LR: to decay or not to decay? Yes, cosine decay with peak LR of $1\text{e}{-3}$ Section 3.2.8
Weight ranked splits? Yes, using normalized scores Section 3.3
Static or dynamic parametrization for the contextualizer? Dynamic parametrization Section 3.3
Replace the contextualizer with self-attention? No Section 3.3
Finally, all training and evaluation runs were executed on 208 NVIDIA H200 GPUs with mixed precision (bfloat16) enabled for training. Across the three model sizes and a 100B-token budget, the total wall-clock training time was approximately 80–90 hours under near-ideal parallelization using the above referenced implementations. To minimize sources of randomness and ensure comparability, we disabled torch.compile during all design-choice and sensitivity experiments. For ablations and final training runs, we enabled torch.compile when compatible to accelerate training. We also fixed the random seed to 11 (arbitrarily chosen) for all runs to further reduce variability due to stochastic effects.
3.2 Design Choices
We conducted more than 200 experiments to investigate key design choices. Table 2 summarizes the main findings and cites the corresponding experiments that support each conclusion. We next elaborate on all these experiments.
3.2.1 Activation or No Activation in the Enricher
Table 3: Avey’s performance without and with an activation function in the enricher. The study isolates the neural processor (excluding the ranker) and trains it on 10B tokens.